# NeB-SLAM: Neural Blocks-based Salable RGB-D SLAM for Unknown Scenes

### v1

@@ Abstract
Neural implicit representations have recently demonstrated considerable potential in the field of visual simultaneous localization and mapping (SLAM). This is due to their inherent advantages, including low storage overhead and representation continuity. However, these methods necessitate the size of the scene as input, which is impractical for unknown scenes. Consequently, we propose NeB-SLAM, a neural block-based scalable RGB-D SLAM for unknown scenes. Specifically, we first propose a divide-and-conquer mapping strategy that represents the entire unknown scene as a set of sub-maps. These sub-maps are a set of neural blocks of fixed size. Then, we introduce an adaptive map growth strategy to achieve adaptive allocation of neural blocks during camera tracking and gradually cover the whole unknown scene. Finally, extensive evaluations on various datasets demonstrate that our method is competitive in both mapping and tracking when targeting unknown environments.
@@ I Introduction
To realize neural dense visual SLAM that is scalable to unknown scenes, we propose NeB-SLAM, an end-to-end neural RGB-D visual SLAM system. Our approach revolves around a divide-and-conquer mapping strategy and an adaptive growth strategy for mapping as illustrated in Fig. 1. To analyze an unfamiliar scene, we begin by initializing a neural block (NeB), which is a cube with a fixed size, based on the current camera pose. Our method then adaptively allocates new NeBs as RGB-D image sequences are input, eventually partitioning the entire unknown environment into multiple submaps, also known as NeBs. In our method, each NeB represents the local scene as a multi-resolution hash grid[17], which allows for high convergence speed and the representation of high-frequency local features. Inspired by Co-SLAM[14], each NeB also encodes the coordinates using One-blob[18] to encourage surface coherence. We conducted a comprehensive assessment of numerous indoor RGB-D sequences and showcased the scalability and predictive capabilities of our approach in unfamiliar settings. In summary, our contributions are as follows:
@@ III-C Neural Block Allocation
Comp. Ratio[<5cm%$\uparrow$ 94.35 95.2 93.72 96.36 94.48 91.87 91.15 90.95 93.51 95.21 67.45 91.90 94.15 87.03 86.82 83.61 86.60
NeB-SLAM Depth L1[cm]$\downarrow$ 0.97 0.77 2.31 1.22 1.39 1.75 1.58 1.47 1.43 3.21 5.72 2.06 1.24 1.37 4.47 3.40 3.07
Acc.[cm]$\downarrow$ 1.99 1.57 1.89 1.52 1.26 2.70 3.23 2.28 1.97 2.06 4.29 1.69 1.88 1.59 2.73 7.04 3.04
Comp.[cm]$\downarrow$ 2.12 1.76 1.95 1.53 1.62 2.39 2.71 2.48 2.07 2.15 6.01 3.07 2.26 2.68 2.48 3.15 3.11
Comp. Ratio[<5cm%$\uparrow$ 94.59 95.46 93.56 96.32 94.62 92.10 91.06 91.18 93.61 94.52 66.63 89.80 94.56 86.92 87.90 84.06 86.34
NeB-SLAM Depth L1[cm]$\downarrow$ 1.02 0.72 2.33 1.18 1.43 2.09 1.56 1.44 1.47 3.38 6.13 1.99 1.31 1.38 4.50 3.48 3.17
Acc.[cm]$\downarrow$ 2.06 1.60 1.93 1.53 1.30 2.74 2.87 2.27 2.04 2.10 4.35 1.87 2.02 1.55 2.79 5.47 2.88
Comp.[cm]$\downarrow$ 2.06 1.75 1.94 1.55 1.58 2.42 2.72 2.52 2.06 2.25 5.21 3.53 2.30 2.62 2.57 3.78 3.18
Comp. Ratio[<5cm%$\uparrow$ 94.79 95.39 93.52 96.27 94.84 91.12 91.6 90.98 93.56 94.57 67.19 89.89 94.23 87.29 87.27 83.66 86.30
@@ IV-A3 Baselines
The present study compares the reconstruction quality and camera tracking of three different methods: iMAP[11], NICE-SLAM[12], and Co-SLAM[14]. iMAP is the reimplemention released by the authors of NICE-SLAM. Prior to the comparison, all methods implement the mesh culling strategy previously described. It is important to note that these methods are designed for known scenes, whereas our method is intended for unknown scenes. This distinction is evident in the input parameters of our method, which do not include the scene size, in contrast to the aforementioned methods.
@@ IV-A4 Implementation Details
These methods are executed on a desktop PC with an Intel Core i9-14900KF CPU and NVIDIA RTX 4090 GPU. In our method (NeB-SLAM), during camera tracking, a sample of $N_{t}=1024$ pixels is taken and 10 iterations are performed to optimize the camera pose. During mapping, we sample $N_{m}=2048$ and $N_{a}=512$ pixels, and utilize 200 iterations for the first frame mapping and 10 iterations for each subsequent five frames of BA. For each ray, we uniformly sample 32 points and depth-guided sample 11 points. For each NeB ($5\times 5\times 5$ m3), a 16-level HashGrid from Rmin = 16 to Rmax is employed, with a maximum hash entries $T=2^{15}$ per level, Rmax is determined by the voxel size of 2 cm, and OneBlob encodes 16 bins per dimension. Additionally, we provide smaller memory versions (NeB-SLAM) that utilize an $T=2^{14}$ HashGrid for each NeB. Two 2-layer shallow MLPs with 32 neurons are utilized to decode SDFs and colors. The geometric feature $\bm{g}$ has a size of 15. We optimize the camera pose using a learning rate of $31e-$ 3 in tracking and the feature grids, decoders, and camera poses during mapping using learning rates of $21e-$ 2, $21e-$ 2 and $31e-$ 3. The weights for each loss are $\lambda_{c}=5$, $\lambda_{d}=0.1$, $\lambda_{sdf}=1000$, $\lambda_{fs}=10$, and $\lambda_{reg}=1e-6$. The truncation distance, $tr$, is set to 10 cm. The learning rate of camera pose in tracking for TUM dataset is set to $21e-$ 2 and we set $tr$ and iterations of BA to 5 cm and 20 respectively. For ScanNet, we uniformly sample 96 points and depth-guided sample 21 points.
@@ IV-B1 Replica dataset
Tab I presents a comparative analysis of the reconstruction accuracy of the proposed method (NeB-SLAM) with that of the baseline method on Replica dataset[56]. Our method achieves comparable outcomes to those observed in Co-SLAM, with the exception of a few scenarios where our approach demonstrates superior performance. Furthermore, our method exhibits more favorable average values for each metric within the dataset. It is observed that the less memory version (NeB-SLAM) yields better results in certain instances. It is noteworthy that our approach does not require the size of the scene to be inputted, and instead relies on the proposed adaptive map growth strategy to gradually allocate NeBs to cover the entire scene. Consequently, in terms of memory, our method is not dominant in some scenes, as shown in Tab. VI. On this dataset, our method allocates 3.7 NeBs equally for a single scene. Since each NeB represents only a localized scene, as illustrated in Fig. 1, our method is more expressive than Co-SLAM, which employs a single hash-grid to represent the entire scene. We also evaluated the absolute trajectory error of our method On this dataset, which is also superior to other baseline methods as shown in Tab. II.
@@ IV-B3 ScanNet and TUM datasets
Following [12] and [14], we also evaluate the tracking accuracy of our method on ScanNet[57] and TUM[59] datasets. The results are presented in Tab. IV and Tab. V. For ScanNet dataset, 6 scenes are evaluated and compared with the baseline methods, resulting in the highest tracking accuracy in most scenes, with an average tracking frame rate of approximately 15.8 FPS. On this dataset, our method allocates an average of 5 NeBs per scene, which is less advantageous in terms of the number of parameters compared to Co-SLAM[14]. However, when targeting unknown scenes, exploring the scene using a small number of redundant parameters is unavoidable. Fig. 5 shows the qualitative comparison on ScanNet scene0000 with different shading mode. Our methods achieve accurate scene reconstruction without the need for scene size. For the TUM dataset, we evaluate 3 of the scenes with an average tracking frame rate of approximately 17.2 FPS. On average, 1.6 NeBs are allocated to each scene, and the number of parameters is less than that of Co-SLAM.
@@ IV-C Performance Analysis
On a desktop PC with an Intel Core i9-14900KF CPU and NVIDIA RTX 4090 GPU, our method (NeB-SLAM) achieves a tracking frame rate of 20 fps when utilising the default settings. For datasets that present greater challenges, such as those from Scannet and TUM, 15 FPS can still be achieved as shown in Tab. VI. In comparison to Co-SLAM, our method does not exhibit superior processing efficiency or a smaller number of parameters. However, we are able to construct a dense map of unknown environments with a minimal increase in parameters, a feat that is not achievable by all the baseline methods. Furthermore, the computational complexity of our method remains relatively constant as the number of NeBs increases, a consequence of our approach, which considers only the NeBs within the current view frustum.
@@ IV-D Ablation Study
In our approach, we primarily consider the impact of hash table size $T$ on the method. We evaluate two sizes of hash tables, as shown in Tab. VI, NeB-SLAM ($T$=14) is more advantageous than NeB-SLAM ($T$=15) in terms of processing efficiency and the number of parameters in each dataset. However, NeB-SLAM yields superior results in terms of reconstruction accuracy and pose estimation. A larger hash table size was not tested since the number of parameters would be significantly higher.
@@ V Conclusion
The proposed NeB-SLAM is designed to address the challenge of constructing dense maps for unknown scenes. Our approach involves a divide-and-conquer strategy, whereby the unknown scene is divided into multiple NeBs of fixed size. These NeBs are adaptively allocated during camera tracking, gradually covering the entire unknown scene. Furthermore, when confronted with larger scenes, our method ensures the linear growth of model parameters, rather than geometric growth, while maintaining the scene representation capability of the model.

### v2

@@ Abstract
Neural implicit representations have recently demonstrated considerable potential in the field of visual simultaneous localization and mapping (SLAM). This is due to their inherent advantages, including low storage overhead and representation continuity. However, these methods necessitate the size of the scene as input, which is impractical for unknown scenes. Consequently, we propose NeB-SLAM, a neural block-based scalable RGB-D SLAM for unknown scenes. Specifically, we first propose a divide-and-conquer mapping strategy that represents the entire unknown scene as a set of sub-maps. These sub-maps are a set of neural blocks of fixed size. Then, we introduce an adaptive map growth strategy to achieve adaptive allocation of neural blocks during camera tracking and gradually cover the whole unknown scene. Furthermore, the cumulative drift is corrected through global loop closure detection and global Bundle Adjustment. Finally, extensive evaluations on various datasets demonstrate that our method is competitive in both mapping and tracking when targeting unknown environments.
@@ I Introduction
To realize neural dense visual SLAM that is scalable to unknown scenes, we propose NeB-SLAM, an end-to-end neural RGB-D visual SLAM system. Our approach revolves around a divide-and-conquer mapping strategy and an adaptive growth strategy for mapping as illustrated in Fig. 1. To analyze an unfamiliar scene, we begin by initializing a neural block (NeB), which is a cube with a fixed size, based on the current camera pose. Our method then adaptively allocates new NeBs as RGB-D image sequences are input, eventually partitioning the entire unknown environment into multiple submaps, also known as NeBs. In order to identify the global position of each input frame, we employ the use of the Bag of Words (BoW)[17] for the purpose of online detection of loop closure. Upon the occurrence of a loop closure, we proceed to undertake a global Bundle Adjustment (BA) in order to rectify the cumulative drift that has occurred between the trajectories of the closed loops. In our method, each NeB represents the local scene as a multi-resolution hash grid[18], which allows for high convergence speed and the representation of high-frequency local features. Inspired by Co-SLAM[14], each NeB also encodes the coordinates using One-blob[19] to encourage surface coherence. We conducted a comprehensive assessment of numerous indoor RGB-D sequences and showcased the scalability and predictive capabilities of our approach in unfamiliar settings. In summary, our contributions are as follows:
@@ III-C Neural Block Allocation
Comp. Ratio[<5cm%$\uparrow$ 94.35 95.20 93.72 96.36 94.48 91.87 91.15 90.95 93.51 95.21 67.45 91.90 94.15 87.03 86.82 83.61 86.60
NeB-SLAM Depth L1[cm]$\downarrow$ 0.95 0.75 2.25 1.12 1.42 1.73 1.35 1.44 1.38 3.13 5.02 2.06 1.21 1.28 4.04 2.89 2.80
Acc.[cm]$\downarrow$ 1.85 1.54 1.76 1.45 1.31 2.56 2.59 2.28 1.92 1.87 4.11 1.69 1.76 1.55 2.77 6.24 2.86
Comp.[cm]$\downarrow$ 2.03 1.76 1.57 1.36 1.67 2.31 2.37 2.48 1.94 1.88 5.35 3.07 2.20 2.51 2.37 3.25 2.95
Comp. Ratio[<5cm%$\uparrow$ 94.78 95.56 93.89 97.32 94.67 92.15 91.24 91.18 93.85 95.52 67.63 89.80 94.67 86.98 87.92 84.13 86.66
NeB-SLAM Depth L1[cm]$\downarrow$ 1.03 0.77 2.34 1.18 1.45 1.89 1.55 1.44 1.46 3.27 5.13 1.99 1.33 1.33 4.38 3.12 2.94
Acc.[cm]$\downarrow$ 1.98 1.60 1.91 1.55 1.37 2.69 2.65 2.27 2.00 2.02 4.56 1.87 2.11 1.51 2.77 5.47 2.90
Comp.[cm]$\downarrow$ 2.05 1.76 1.94 1.55 1.58 2.45 2.70 2.52 2.07 2.16 5.09 3.53 2.31 2.61 2.46 3.78 3.13
Comp. Ratio[<5cm%$\uparrow$ 94.79 95.39 93.40 96.78 94.84 91.67 91.21 90.98 93.63 95.13 67.33 89.89 94.22 87.18 87.33 83.66 86.39
@@ III-D Tracking and Mapping
III-E Loop Closure
To address arbitrary drift, a BoW[17] model for global position identification is utilized, wherein each global keyframe is incorporated. Upon the generation of a global keyframe, it is appended to the aforementioned database. This methodology contrasts with that of MIPSFusion[58], which employs submap overlap for the detection of loop closures and is constrained to the correction of smaller drifts.
For each new keyframe with camera pose $\bm{T}_{c}=[\bm{R}|\bm{t}]$, a loop closure is identified by querying the BoW database. The first $K$ keyframes from BoW that are not adjacent to the current keyframe are queried, and the two frames $\bm{T}_{r1}$ and $\bm{T}_{r2}$ with the highest visual similarity score and the greatest timestamp distance are identified as the closed-loop keyframes. The threshold is defined as the minimum similarity score between the current keyframe and its neighboring keyframes. For each closed-loop keyframe, the reprojection error between it and the current keyframe is calculated in order to optimize the pose of the latter:
$\begin{split}&\bm{T}_{c}=\arg\min_{\bm{T}_{c}}\sum_{i=1}^{2}\sum_{j=1}^{N}\|\bm{u}_{ij}^{\prime}-\bm{u}_{j}\|_{2}^{2}.\end{split}$ (14)
The process commences with the back-projection of the pixels $\bm{u}_{j}$ belonging to the current keyframe into the world coordinate system $\bm{P}_{j}=\bm{T}_{c}\pi^{-1}(\bm{u}_{j})$. Subsequently, the point is projected onto the corresponding closed-loop keyframe $\bm{u}_{ij}=\pi(\bm{T}_{ri}^{-1}\bm{P}_{j})$, and the corresponding depth value is obtained by interpolating the corresponding pixel coordinates with the depth map. Subsequently, the depth value is utilized to backproject $\bm{u}_{ij}$ to the world coordinate system and project it to the current keyframe $\bm{u}_{ij}^{\prime}=\pi(\bm{T}_{c}^{-1}\bm{T}_{ri}\pi^{\prime-1}(\bm{u}_{ij}))$. Here, $\pi$ is used to perform dehomogenization and perspective projection and $\pi^{\prime}$ employs the depth data obtained through interpolation. Subsequently, the optimized current pose $\bm{T}_{c}$ and relative poses between frames are employed to adjust the keyframe poses between closed loops. Ultimately, the camera poses and map between the closed loops are optimized in two stages using the aforementioned method in Sec. III-D. Initially, the camera poses are fixed in order to optimize the map. Thereafter, the camera poses and maps are optimized concurrently.
@@ IV-A3 Baselines
The present study compares the reconstruction quality and camera tracking of the following different methods: iMAP[11], NICE-SLAM[12], Co-SLAM[14], Vox-Fusion[62], MIPS-Fusion[58] and SplaTAM[63]. iMAP is the reimplemention released by the authors of NICE-SLAM. Prior to the comparison, all methods implement the mesh culling strategy previously described. It is important to note that these methods are designed for known scenes, whereas our method is intended for unknown scenes. This distinction is evident in the input parameters of our method, which do not include the scene size, in contrast to the aforementioned methods.
@@ IV-A4 Implementation Details
These methods are executed on a desktop PC with an Intel Core i9-14900KF CPU and NVIDIA RTX 4090 GPU. In our method (NeB-SLAM), $\tau_{th}$ is set to 0.2. During camera tracking, a sample of $N_{t}=1024$ pixels is taken and 10 iterations are performed to optimize the camera pose. During mapping, we sample $N_{m}=2048$ and $N_{a}=512$ pixels, and utilize 200 iterations for the first frame mapping and 10 iterations for each subsequent five frames of BA. For each ray, we uniformly sample 32 points and depth-guided sample 11 points. For each NeB ($5\times 5\times 5$ m3), a 16-level HashGrid from Rmin = 16 to Rmax is employed, with a maximum hash entries $T=2^{15}$ per level, Rmax is determined by the voxel size of 2 cm, and OneBlob encodes 16 bins per dimension. Additionally, we provide smaller memory versions (NeB-SLAM) that utilize an $T=2^{14}$ HashGrid for each NeB. Two 2-layer shallow MLPs with 32 neurons are utilized to decode SDFs and colors. The geometric feature $\bm{g}$ has a size of 15. We optimize the camera pose using a learning rate of $31e-$ 3 in tracking and the feature grids, decoders, and camera poses during mapping using learning rates of $21e-$ 2, $21e-$ 2 and $31e-$ 3. The weights for each loss are $\lambda_{c}=5$, $\lambda_{d}=0.1$, $\lambda_{sdf}=1000$, $\lambda_{fs}=10$, and $\lambda_{reg}=1e-6$. The truncation distance, $tr$, is set to 10 cm. The learning rate of camera pose in tracking for TUM dataset is set to $21e-$ 2 and we set $tr$ and iterations of BA to 5 cm and 20 respectively. For ScanNet, we uniformly sample 96 points and depth-guided sample 21 points.
@@ IV-B1 Replica dataset
Tab I presents a comparative analysis of the reconstruction accuracy of the proposed method (NeB-SLAM) with that of the baseline method on Replica dataset[57]. In the majority of instances, our method demonstrates superior performance. In comparison to Co-SLAM, our method has demonstrated enhanced performance in nearly all scenarios. It is observed that the less memory version (NeB-SLAM) yields better results in certain instances. It is noteworthy that our approach does not require the size of the scene to be inputted, and instead relies on the proposed adaptive map growth strategy to gradually allocate NeBs to cover the entire scene. Consequently, in terms of memory, our method is not dominant in some scenes, as shown in Tab. VI. On this dataset, our method allocates 3.7 NeBs equally for a single scene. Since each NeB represents only a localized scene, as illustrated in Fig. 1, our method is more expressive than Co-SLAM[14], which employs a single hash-grid to represent the entire scene. Furthermore, the absolute trajectory error of our method is evaluated on this dataset, as illustrated in Tab. II. Our method is demonstrated to outperform other baseline methods, with the exception of the 3D Gaussian Splatting-based SplaTAM[63].
@@ IV-B3 ScanNet and TUM datasets
Following [12] and [14], we also evaluate the tracking accuracy of our method on ScanNet[59] and TUM[61] datasets. The results are presented in Tab. IV and Tab. V. For ScanNet dataset, 6 scenes are evaluated and compared with the baseline methods, resulting in the highest tracking accuracy in most scenes, with an average tracking frame rate of approximately 15.8 FPS. On this dataset, our method allocates an average of 5 NeBs per scene, which is less advantageous in terms of the number of parameters compared to Co-SLAM[14]. However, when targeting unknown scenes, exploring the scene using a small number of redundant parameters is unavoidable. Fig. 5 shows the qualitative comparison on ScanNet scene0000 with different shading mode. Our methods achieve accurate scene reconstruction without the need for scene size. For the TUM dataset, we evaluate 3 of the scenes with an average tracking frame rate of approximately 17.2 FPS. On average, 1.6 NeBs are allocated to each scene, and the number of parameters is less than that of Co-SLAM. It is noteworthy that both datasets are based on real-world scenes. The tracking accuracy of our method on both datasets is superior to that of SplaTAM. Moreover, both Vox-Fusion[62] and MIPS-Fusion[58] can be employed in situations where the specific circumstances are not yet known. However, Vox-Fusion lacks the functionality of loop closure detection, which impedes the correction of cumulative drift. MIPS-Fusion employs a submap overlap approach to detect loop closure, which is only capable of correcting smaller drifts.
@@ IV-C Performance Analysis
On a desktop PC with an Intel Core i9-14900KF CPU and NVIDIA RTX 4090 GPU, our method (NeB-SLAM) achieves a tracking frame rate of 20 fps when utilising the default settings. For datasets that present greater challenges, such as those from Scannet and TUM, 15 FPS can still be achieved as shown in Tab. VI. In comparison to Co-SLAM, our method does not exhibit superior processing efficiency or a smaller number of parameters. Nevertheless, we are capable of constructing a comprehensive map of uncharted environments with a minimal increase in parameters, while ensuring high-precision tracking. This is a capability that is not available with all baseline methods. Furthermore, the computational complexity of our method remains relatively constant as the number of NeBs increases, a consequence of our approach, which considers only the NeBs within the current view frustum.
TABLE VII: Ablation study on loop closure on Replica dataset.
[…]
0.84 1.43 1.97 2.07 93.61
0.59 1.38 1.92 1.94 93.85
@@ IV-D Ablation Study
We evaluate two sizes of hash tables, as shown in Tab. VI, NeB-SLAM ($T$=14) is more advantageous than NeB-SLAM ($T$=15) in terms of processing efficiency and the number of parameters in each dataset. However, NeB-SLAM yields superior results in terms of reconstruction accuracy and pose estimation. A larger hash table size was not tested since the number of parameters would be significantly higher.
Furthermore, the impact of loop closure detection on tracking accuracy and reconstruction quality is evaluated on Replica dataset[57], as illustrated in Tab. VII. The implementation of global pose correction through the use of loop closure detection has been demonstrated to enhance the precision of camera tracking. The attainment of high-accuracy camera poses has been shown to facilitate greater global consistency in the mapping process, which in turn leads to an improvement in the quality of the reconstructed map.
@@ V Conclusion
The proposed NeB-SLAM is designed to address the challenge of constructing dense maps for unknown scenes. Our approach involves a divide-and-conquer strategy, whereby the unknown scene is divided into multiple NeBs of fixed size. These NeBs are adaptively allocated during camera tracking, gradually covering the entire unknown scene. The BoW model is also employed for global loop closure detection with the objective of correcting the cumulative error. This results in enhanced camera tracking accuracy and global map consistency. Furthermore, when confronted with larger scenes, our method ensures the linear growth of model parameters, rather than geometric growth, while maintaining the scene representation capability of the model.